The random variable Y has the distribution function

F(y)= 0 y<1

ln 0<y<1

1y>1

find the following, i) P( y>2) ii) P(2<y<2)


1
Expert's answer
2021-04-19T04:53:04-0400

"F_Y(y) = \begin{cases}
 0 &\text{if } y\leq0 \\
 ln\ y &\text{if } 0< y\leq1 \\
 1 &\text{if } y>1 \\
\end{cases}"


i) "P(Y>2)= 1-P(Y\leq2)=1-F_Y(2)=0"


ii) "P(1<Y<2)=F_Y(2)-F_Y(1)=1-ln1=1"